# Machine-Learning-Enhanced Optimization of Noise-Resilient Variational Quantum Eigensolvers

### v1

@@ 4.1 Hardware Noise Analysis of EMICoRe
Previous work [9] analyzed EMICoRe’s performance when measurement shot noise is present. It was observed, that EMICoRe outperforms NFT more significantly for higher levels of measurement shot noise, i.e., less shots. In this proceeding, we investigate the algorithm’s performance when accounting for hardware noise as well.
In the following, the setup is kept constant for all experiments with $5$-qubits and $3$-layers of the EfficientSU2 circuit and the Ising Hamiltonian with open boundary conditions. All simulations are performed for a total of $600$ measurements  [^3] Note that two measurements are performed for each step, i.e., $600$ measurements correspond to $300$ optimization steps. and $N=1024$ shots per measurement are used.
In Fig. 3(a) we run experiments with simulated hardware noise and no error mitigation technique while in Fig. 3(b) and Fig. 3(c) we use twisted readout error extinction (TREX) [19] and zero noise extrapolation (ZNE) [20, 21] as error mitigation strategies. The hardware noise is simulated with the Fake5QV1 backend. The results in Fig. 3(a) are obtained by aggregating $50$ independent runs, while Figs. 3(b) and 3(c) uses only $10$ runs, as the feature was deprecated before more data could be aggregated. Table 1 summarizes the results from Fig. 3, with mean and standard deviations computed over multiple seeded trials, for simulations with and without hardware noise.
When hardware noise is considered, we observe that EMICoRe significantly outperforms NFT, see Fig. 3(a), converging faster and more reliably than NFT, e.g., the trial density on the right-hand-side is more peaked around the median. We empirically observe that NFT struggles significantly more with hardware noise compared to EMICoRe. Notably, neither algorithm achieves the ground-state energy; however, EMICoRe attains a fidelity of nearly $80\%$. When using error mitigation, EMICoRe outperforms NFT in both energy minimization and fidelity, being more stable with respect to different initializations, and achieving both lower energies and higher fidelities.
Table 1: Mean and standard deviation of energy and fidelity calculated at the last optimization step, for all simulations shown in Fig. 3. For the energy lower is better, for the fidelity higher is better.
@@
Energy $f^{\ast}({\boldsymbol{\hat{\theta}}})$ Fidelity Energy $f^{\ast}({\boldsymbol{\hat{\theta}}})$ Fidelity
@@ 5 Conclusion
In these proceedings, we leveraged a novel machine-learning-enhanced optimization of VQE algorithms for finding the ground state of the quantum Ising Hamiltonian in the presence of simulated hardware noise. This algorithm, termed EMICoRe [9], builds on Gaussian Process Regression, for noise-resilient estimation of the energy, and Bayesian Optimization, to predict the new candidate points to measure at different step in the optimization.
Our numerical experiments validated the capability of EMICoRe to outperform the state-of-the-art baseline, the NFT algorithm [7], in the presence of simulated hardware noise. Even in the presence of simulated hardware noise, EMICoRe is able to converge to lower energies compared to NFT and is more robust with respect to different initializations. This robustness highlights the potential of the proposed method for future applications to more complicated quantum systems, and under the challenging conditions of real NISQ devices. As a next step, we aim to scale up EMICoRe for larger systems and ultimately evaluate its performance on real quantum hardware.

### v2

@@ 4.1 Hardware Noise Analysis of EMICoRe
Previous work [9] analyzed EMICoRe’s performance in the presence of measurement shot noise, showing that EMICoRe outperforms NFT more significantly at higher noise levels, i.e., with fewer shots. In these proceedings, we extend the analysis to include hardware noise as well.
In the following experiments, the setup remains constant: we consider a $5$-qubit system with $3$-layers of the EfficientSU2 circuit and the Ising Hamiltonian with open boundary conditions. All simulations are conducted with a total of $600$ measurements  [^3] Note that two measurements are performed at each step, meaning that $600$ measurements correspond to $300$ optimization steps., using $N=1024$ shots per measurement.
In Fig. 3(a), we conduct experiments with simulated hardware noise and no error mitigation. In Fig. 3(b) and Fig. 3(c), we apply twisted readout error extinction (TREX) [18] and zero noise extrapolation (ZNE) [19, 20] as error mitigation strategies, respectively. Hardware noise is simulated with the Fake5QV1 backend. The results in Fig. 3(a) are obtained by aggregating $50$ independent runs, whereas Figs. 3(b) and 3(c) include only $10$ runs, as the feature was deprecated before further data could be collected. Table 1 summarizes the results from Fig. 3, reporting the mean and standard deviation computed over multiple seeded trials for simulations with and without error mitigation.
When hardware noise is taken into account, EMICoRe significantly outperforms NFT, as shown in Fig. 3(a), demonstrating faster and more reliable convergence. For instance, the trial density on the right-hand-side is more peaked around the median. These empirical results indicate that NFT is considerably more affected by hardware noise compared to EMICoRe. While neither algorithm reaches the ground-state energy, EMICoRe achieves a fidelity of nearly $80\%$. With error mitigation, EMICoRe outperforms NFT in both energy minimization and fidelity, demonstrating greater stability across different initializations and achieving lower energies as well as higher fidelities.
Table 1: Mean and standard deviation of the energy and fidelity at the final optimization step for all simulations presented in Fig. 3. For the energy, lower values are better; for the fidelity, higher values are better.
@@
Energy $E^{\ast}({\boldsymbol{\hat{\theta}}})$ Fidelity Energy $E^{\ast}({\boldsymbol{\hat{\theta}}})$ Fidelity
@@ 5 Conclusion
In these proceedings, we presented a novel machine-learning-enhanced optimization approach for VQE algorithms to find the ground state of the quantum Ising Hamiltonian in the presence of simulated hardware noise. This algorithm, called EMICoRe [9], combines Gaussian Process Regression for noise-resilient energy estimation and Bayesian Optimization to predict new candidate points for measurement at each step of the optimization.
Our numerical experiments demonstrated that EMICoRe outperforms the state-of-the-art baseline, the NFT algorithm [7], in the presence of simulated hardware noise. EMICoRe converges to lower energies than NFT and shows greater robustness across different initializations. This robustness underscores the potential of the proposed method for future applications to more complex quantum systems and under the challenging conditions of real NISQ devices.